OpenStudy (anonymous):

solve for x: 1. log6(x+3)+log6(x+4) = 1 "log base 6"

OpenStudy (anonymous):

hmm why divide?

OpenStudy (anonymous):

log M + log N = Log MN right?

myininaya (myininaya):

\[\log_6[(x+3)(x+4)]=1\] \[(x+3)(x+4)=6\] \[x^2+4x+3x+12=6\] \[x^2+7x+12-6=0\] \[x^2+7x+6=0\] \[(x+6)(x+1)=0\] \[x=-6,x=-1\] but only x=-1 works

OpenStudy (anonymous):

-6 doesnt work because it makes the log negative right?

myininaya (myininaya):

yes -6+4 is negative -6+3 is negative